OpenStudy (anonymous):

y varies directly as the square of r. If y=162 when r=9, find y when r=5. Y=?

OpenStudy (anonymous):

50

OpenStudy (ash2326):

y varies directly withsquare of r y=kr^2 k is a constant y=162 when r=9 162=k*81 k=2 y=2r^2 r=5 y=2*5^2 y=50

OpenStudy (king):

50